The length of a rectangle is 5 ft longer than twice the width. if the perimeter is 100ft, find the length and width of the rectangle.
Answer by Roseghanezadeh(16)

Let's give the width the variable "w" and the length the variable "l". So it says that the length of a rectangle is 5 ft longer than twice the width which we can write as "5 + 2w". And that its perimeter = 100ft.
We know the formula for the perimeter of a rectangle is:
2l + 2w = p
Now we substitute:
2(5 + 2w) + 2w = 100
10 + 4w + 2w = 100
Now we take like terms to one side:
6w = 90
w = 15ft
We also knew that "l = 5 + 2w", so:
l = 5 + 2(15)
l = 35ft
Check with the formula "2l + 2w = p"
2(35) + 2(15) = 100
70 + 30 = 100
100 = 100